solve for the maximum torque within uniform electric field of 6 x 10^4N/C experiences a dipole moment of 5 x 10^-14Cm.


1
Expert's answer
2022-02-20T15:47:41-0500

Answer

maximum torque within uniform electric field of 6 x 10^4N/C experiences a dipole moment of 5 x 10^-14Cm

"\tau=PEsin\theta\\=5*10^{-14}*6*10^4 *sin90°\\=30*10^{-10}Nm"